The bearing of a lighthouse from ship A is 29°NE and the bearing of that same
lighthouse from ship B is 61°NW. If ship A and ship B are on the east-west line 200 km
apart, how far is ship B from the lighthouse?
1
Expert's answer
2020-09-27T14:04:01-0400


y - distance between lighthouse and ship B

"61\degree NW + 29\degree NE = 90\degree WE"

ABC is a right triangle.

"\cos\angle BAC = \dfrac{AB}{AC}"

"\cos(90\degree-61\degree)=\dfrac{y}{200}"

"\cos(29\degree)=\dfrac{y}{200}"

"y=\cos(29\degree)*200"

"y=175" km

Answer: "175" km is ship B from the lighthouse.